### v1

# Simulating Prosocial Behavior and Social Contagion in LLM Agents under Institutional Interventions

Abstract
As large language models (LLMs) increasingly serve as autonomous agents in social contexts, understanding their capacity for prosocial behavior becomes essential. We present ProSim, a simulation framework designed to examine how prosocial behavior emerges, adapts, and erodes in LLM-based agents under diverse social and institutional conditions. The framework comprises four components: individual simulation, scenario simulation, interaction simulation, and intervention simulation. We conduct three progressive studies to evaluate prosocial alignment. First, we show that LLM agents can demonstrate stable and context-sensitive prosocial behavior across diverse scenarios and adapt their responses under normative policy interventions. Second, we find that agents engage in fairness-based third-party punishment and respond systematically to variations in inequity magnitude and enforcement cost. Third, we show that policy-induced inequities suppress prosocial behavior, propagate through social networks, and are mediated by agents’ perceptions of unfairness. These findings lay the groundwork for evaluating social alignment and modeling institutional dynamics in agent-driven societies.
1 Introduction
Large language models (LLMs) have evolved beyond next-word prediction into general-purpose agents equipped with complex reasoning, decision-making, and social interaction capabilities zhao2023survey ; gpt4 . Recent studies suggest that LLMs can exhibit emergent social-cognitive abilities mou2024individual ; piao2025agentsociety , including theory of mind li2023theory ; strachan2024testing , moral judgment ramezani2023knowledge , and value alignment liu2022aligning . These developments position LLMs as promising tools for simulating human-like behavior in synthetic populations.
This work focuses on prosocial behavior, defined as voluntary actions intended to benefit others or promote collective welfare penner2005prosocial ; grueneisen2022development . Prosocial behavior plays a critical role in sustaining cooperation, trust, and social cohesion cameron2022empathy ; carlo2005interplay ; thielmann2020personality . Prior research brookins2023playing ; xie2024can has primarily examined prosociality in LLMs through economic game paradigms such as the dictator game and public goods game. However, existing studies face two key limitations. First, they narrowly frame prosociality in terms of economic cooperation (e.g., resource allocation), overlooking diverse real-world expressions such as donating and volunteering. Second, they rely on static, single-shot evaluations, neglecting how prosocial behavior dynamically evolves in changing environments or in response to institutional interventions.
To address these limitations, we propose ProSim, a comprehensive simulation framework that models the emergence and evolution of prosocial behavior in LLM-based agents. The framework consists of four components. (1) Individual simulation module, which instantiates each agent with demographic attributes and psychological traits such as empathy and moral identity. (2) interaction simulation module, which places agents in a small-world network and supports repeated multi-agent interactions. (3) Scenario simulation module, which reproduces six distinct prosocial tasks: helping, donating, volunteering, cooperating, information sharing, and recycling. (4) Intervention simulation module, which implements prosocial policy interventions and allows manipulation of fairness conditions through reward asymmetry and burden asymmetry. Building on this framework, we conduct a series of simulation studies to investigate the social capacities of LLM agents from multiple perspectives.
We begin by evaluating whether LLM agents naturally exhibit prosocial behavior in structured contexts and how they adjust when exposed to policy-based interventions. Drawing from established behavioral paradigms, we design six typical prosocial scenarios and observe the agents’ default tendencies. To assess their responsiveness to external cues, we introduce four types of prosocial policy interventions and evaluate how these modulate agent behavior. Our results show that LLMs agents can exhibit stable and context-sensitive prosocial behavior across diverse scenarios, and adjust their behavior in response to policy interventions.
In the second study, we test whether LLM agents can perceive and respond to inequity by enforcing social norms. We adapt a third-party punishment paradigm fehr2004third in which agents observe unfair resource distributions and must decide whether to penalize the transgressor at a personal cost. In addition to behavioral responses, we analyze agents’ emotional expressions to assess their affective alignment with human fairness reasoning. Using a comparable human dataset for benchmarking, we find that LLM agents are capable of norm-enforcing third-party punishment, showing sensitivity to both the degree of unfairness and the cost of enforcement.
Finally, we explore how prosocial behavior evolves over time under policy-induced inequity in networked environments. Agents are embedded in a small-world social network and repeatedly interact across rounds of simulated exchanges. Two types of structural unfairness (reward asymmetry and burden asymmetry) are introduced and allowed to diffuse through the network. We track the resulting behavioral trajectories to assess whether prosociality is sustained or eroded at the population level. The findings reveal that Policy inequities significantly undermine prosocial behavior in LLM agents, an effect amplified through social contagion and mediated by agents’ perceived unfairness.
Together, these findings suggest that LLM agents can serve as scalable proxies for studying the emergence and evolution of prosocial behavior in synthetic societies. Our contributions are threefold:
We propose ProSim, a modular simulation framework for modeling the emergence and evolution of prosocial behavior in LLM agents. The framework integrates four key modules to approximate the complexity of real human social environments.
We show that LLM agents not only exhibit prosocial behavior but also adapt their responses based on fairness cues, aligning with human behavioral benchmarks.
Leveraging the aligned simulation environment, we model how structural policy inequities affect the emergence, decay, and diffusion of prosocial norms across simulated societies.
2 The ProSim Framework
Simulating prosocial behavior poses a fundamental challenge due to its psychological and social complexity. Human prosociality is shaped not only by internal traits such as empathy and moral identity, but also by contextual factors, social influence, and institutional structures.
To address this, we introduce ProSim, a modular simulation framework for studying how prosocial behavior emerges, adapts, and deteriorates in LLM-based agents across varied social and policy environments. Grounded in computational social psychology edelmann2020computational , ProSim supports controlled experimentation at both individual and collective levels. As shown in Figure 1, the framework comprises four modules: (1) Individual Simulation, (2) Scenario Simulation, (3) Interaction Simulation, and (4) Intervention Simulation. These components jointly model the dynamic interplay between psychological traits, situational context, and institutional interventions in shaping agent behavior.
2.1 Individual Simulation
This module defines the foundational identity of each agent by capturing the heterogeneity inherent in human populations. Agents are initialized with realistic demographic and psychological profiles, which serve as stable priors for downstream behavior.
Demographic Attributes. Each agent is assigned demographic features including age, gender, education, income, and employment status. These are sampled from population-level distributions provided by the National Bureau of Statistics (NBS), ensuring structural diversity aligned with real-world sociodemographic patterns.
Psychological Traits. Each agent is further characterized by two sets of psychological traits: (1) core prosocial dispositions eisenberg1999consistency , including empathic concern, moral identity, altruistic tendency, and social responsibility; and (2) the Big Five personality gosling2003very , including openness, conscientiousness, extraversion, agreeableness, and neuroticism. Trait values are drawn from Gaussian distributions calibrated using meta-analytic norms to reflect empirical variability.
All attributes are encoded into the agent’s natural language prompt, grounding each agent in a contextually rich identity. This design ensures that behavior arises not from static rules but from socially and psychologically plausible profiles. More details are documented in Appendix C.1.
2.2 Scenario Simulation
This module defines the situational contexts in which agents are prompted to make prosocial decisions. This component is essential for evaluating how different types of social dilemmas and environmental cues influence prosocial tendencies under controlled conditions. To ensure comprehensive coverage of real-world prosocial behaviors, we define six distinct scenarios that vary across cost level, norm expectation, reward visibility, and dependency level on others.
Dimension | Volunteering | Helping | Donating | Cooperation | Recycling | Sharing |
Cost Level | High | Medium | Medium | Medium | Low | High |
Norm Expectation | High | High | Medium | Medium | High | Medium |
Reward Visibility | Medium | High | High | High | Low | Medium |
Dependency Level | Medium | High | Low | High | Low | Medium |
As shown in Table 1, the six scenarios (e.g. Volunteering, Helping, Donating, Cooperation, Recycling, Sharing) differ across structural dimensions. The classification of each scenario along the four dimensions is guided by expert judgment from social psychologists. Each scenario is delivered via a standardized prompt, designed to reflect realistic decision-making contexts. Following the prompt, agents respond using a consistent 7-point Likert scale to indicate their level of prosocial intention.
2.3 Interaction Simulation
This module captures how prosocial behavior evolves through repeated agent-to-agent interactions within a structured social network. This layer is essential for modeling emergent dynamics that cannot be explained by individual traits or isolated decisions alone. Concretely, we initialize the agent society using a small-world network generated via the Watts–Strogatz algorithm kleinberg2000small . This structure reflects key properties of real-world social systems, including high local clustering and short average path lengths. Each node represents an LLM-based agent, and each edge denotes a potential communication or observation channel between two agents.
At each simulation timestep , a random subset of edges is activated. Each agent observes the latest prosocial decisions made by neighbors , where is the set of agents connected to via active edges at time . By integrating (i) the shared scenario narrative , (ii) its own prior decision , and (iii) the observed actions of its neighbors , the LLM agent then generates its decision on six prosocial scenarios: Through this module, we simulate how prosocial norms and behaviors diffuse or decay over time within a socially structured community of LLM agents.
2.4 Intervention Simulation
This module enables controlled manipulation of policy conditions to investigate how institutional interventions and structural inequalities influence the development of prosocial behavior. While individual traits and social dynamics contribute to behavioral variation, policy environments play a central role in shaping collective behaviors through top-down regulation.
Prosocial Policies. Prosocial policies aim to promote cooperative and altruistic behavior at the societal level by shaping how agents interpret and respond to social dilemmas. We classify interventions along two orthogonal dimensions: (1) Mechanism of Influence: Cognitive () interventions target internal beliefs and perceptions, whereas Behavioral () interventions directly affect observable actions. (2) Mode of Compliance: Voluntary () policies rely on intrinsic motivation and social persuasion, while Compulsory () policies impose rules or sanctions. Based on this taxonomy, we implement four representative interventions:
Moral Indoctrination (,): appeals to internalized moral values to motivate prosociality.
Regulatory Enforcement (,): mandates behavior through institutional rules or penalties.
Social Comparison (,): exposes agents to peer decisions to activate normative pressure.
Economic Incentives (,): introduces rewards or penalties contingent on agent choices.
Each intervention is embedded into the natural language prompt as contextual information, enabling agents to adapt their decisions based on perceived institutional signals.
Policy inequity. Beyond evaluating idealized interventions, we introduce policy inequity to examine the broader social impact of unfair treatment. In real-world systems, rewards and burdens are often distributed asymmetrically across individuals or groups due to structural bias or implementation constraints. To simulate these conditions, we define two types of inequity:
Reward Asymmetry: agents receive unequal recognition or benefits despite contributing equally.
Burden Asymmetry: agents face unequal costs or effort for performing the same prosocial task.
We randomly assign these asymmetries to a subset of agents and allow their effects to propagate through the social network. This setup enables us to assess how inequitable policies influence the emergence and spread of prosocial behavior within agent societies.
2.5 Experimental Settings
To evaluate the capabilities of the ProSim framework, we conduct three progressive studies that examine how LLM agents exhibit, perceive, and respond to prosocial dynamics.
Study 1: Can LLM Agents Exhibit Prosocial Behavior? This study investigates whether LLM agents can demonstrate consistent and context-sensitive prosocial intentions across six scenarios.
Study 2: Do LLM Agents Perceive and Respond to Inequity? This study tests whether agents can detect unequal treatment and adjust their decisions based on the perceived severity of the inequity.
Study 3: How Does Policy Inequity Affect Prosociality? This study explores how policy-induced inequities influence prosocial motivation and how such effects spread through social networks.
Overall Model and Agent Configuration. We evaluate ProSim using five LLMs, including three open-source models (LLaMA-3-8B llama3 , Qwen-2.5-7B qwen2 , DeepSeek-v3 deepseek ) and two proprietary models (GPT-3.5-turbo gpt35 , GPT-4o gpt4 ). LLaMA and Qwen models are deployed on a single NVIDIA A100 GPU for inference, while DeepSeek and the GPT series are accessed via APIs, with the generation temperature is set to 1.0. We initialize 104 agents sampled from real-world distributions, embedded in a small-world network with neighborhood size and rewiring probability . Additional implementation details are provided in each study and Appendix C.
Human Benchmarking. To contextualize LLM behavior, we conduct parallel experiments with 54 human participants recruited online for Study 1 and Study 2. Each participant completed the same tasks presented to the LLM agents, using identical scenario prompts and response formats. Prior to task completion, participants provided informed consent and completed standardized psychological inventories aligned with the trait dimensions used in agent initialization.
3 Can LLM Agents Exhibit Prosocial Behavior?
This section evaluates whether LLM agents can express context-sensitive prosocial behavior in a manner comparable to humans. We first assess their baseline prosocial intentions across diverse scenarios, then investigate how individual psychological traits shape these behaviors. Finally, we examine whether LLMs adapt their responses under explicit prosocial policy interventions.
3.1 Baseline Prosocial Intention in Diverse Scenarios
We begin by assessing the baseline prosocial tendencies of LLM agents across six representative scenarios. Each agent receives a neutral prompt and rates its willingness to perform a prosocial act on a 7-point Likert scale. Figure 2 (left) shows the distribution of prosocial intentions across scenarios. GPT-3.5 exhibits the highest overall prosociality, particularly in tasks involving helping, donating, and cooperation. GPT-4o and DeepSeek-v3 display relative better alignment with human responses across multiple contexts. Qwen-2.5 shows more polarized behavior, responding positively in low-cost or public-good scenarios, but less so in interpersonal settings such as sharing. LLaMA-3 consistently produces the most conservative ratings across all conditions. We also observe that LLMs tend to be more prosocial in scenarios with high normative expectations and less prosocial when the cost of action is high, suggesting a sensitivity to structural dimensions of prosocial tasks. Figure 2 (right) summarizes average prosociality across all tasks. GPT-3.5 ranks highest overall, followed by human participants, with GPT-4o and Qwen-2.5 at comparable levels. DeepSeek-v3 and LLaMA-3 fall lower on the scale. These results suggest that advanced LLMs can express prosocial intentions that approach human-level patterns, although their responses remain sensitive to scenario context and model-specific characteristics. This baseline evaluation sets the stage for further analysis of individual traits and institutional influences on prosocial behavior.
3.2 The Influence of Psychological Traits on Prosociality
To examine how internal traits shape prosocial behavior, we compute SHAP (SHapley Additive exPlanations) values for each agent, quantifying the contribution of individual psychological traits to prosocial decisions across all scenarios. We focus on two models, DeepSeek-v3 and GPT-4o, that most closely align with human behavior in prior analyses, and compare them with human participants.
Figure 3 summarizes the SHAP analysis of each model. We can observe that GPT-4o shows the closest alignment with human trait influence profiles. In both GPT-4o and humans, Altruistic Tendency emerges as the most dominant factor, followed by strong contributions from Empathic Concern. DeepSeek-v3, by contrast, departs more noticeably from the human pattern. It prioritizes Empathic Concern over Altruistic Tendency, and places greater weight on Moral Identity. Human participants exhibit a more balanced and multidimensional trait contribution. While prosocial core traits dominate, cognitive and interpersonal traits such as Extraversion, Openness, and Neuroticism also play notable roles, reflecting the richer and more heterogeneous basis of human moral judgment. These results demonstrate that LLM agents exhibit interpretable and trait-sensitive behavioral patterns, with both convergences and divergences from human profiles.
3.3 Behavioral Shifts Under Prosocial Policy Interventions
We next evaluate whether LLM agents adjust their prosocial behavior when exposed to external policy cues. Each agent completes the same six scenarios from the baseline condition, with prompts modified to include one of four intervention framings: Economic Incentives, Moral Indoctrination, Regulatory Enforcement, or Social Comparison. We compute the relative change in prosocial intention to assess policy responsiveness. Figure 4 summarizes the results. Intervention effectiveness varies across policy types and models. Regulatory Enforcement yields the most consistent improvement across models and scenarios, suggesting that explicit institutional mandates are broadly effective at eliciting compliance. In contrast, Economic Incentives, while highly effective for human participants, produce mixed effects in LLMs. Only GPT-4o and DeepSeek-v3 exhibit reliable prosocial gains across all six scenarios, though the magnitude of their responses remains below human levels, indicating that monetary cues are less salient for LLMs. At the scenario level, interventions are least effective in the Recycling context. In several models, policy framings even lead to marginal declines in prosociality. This may be due to high baseline prosocial intent in this low-cost, norm-driven scenario, where additional interventions could trigger reactance or reduce intrinsic motivation. The right panel of Figure 4 aggregates the number of positive shifts across all policy-scenario combinations. GPT-4o and DeepSeek-v3 exhibit the strongest and most consistent responsiveness, closely matching human participants. GPT-3.5 and LLaMA-3 show moderate adaptability, while Qwen-2.5 lags behind with the fewest positive responses. Taken together, these findings suggest that advanced LLMs exhibit context-sensitive adjustments to social and normative inputs, a capacity that aligns them more closely with human-like social reasoning.
4 Do LLM Agents Perceive and Respond to Inequity?
This section examines whether LLM agents are capable of recognizing unfairness and responding in norm-enforcing ways. To evaluate this capacity, we adapt a third-party punishment paradigm from behavioral economics fehr2004third , a well-established framework for studying fairness judgment and altruistic punishment. Each agent completes a 60-trial task in which two newly assigned virtual players propose an allocation of $30. Player 1 offers $x to Player 2 and retains $[30-x] for themselves. The participants, acting as third-party judges, then chose between:
Accept: Implement the allocation and receive a $10 reward;
Punish: Pay $y to eliminate Player 1’s earnings; Player 2 keeps $x, and the agent receives $10–y.
We record agents’ binary choices (accept = 0, punish = 1) across trials and and compare the results with humans. Emotional responses are also collected but reported separately in Appendix D.
4.1 Third-Party Punishment as a Response to Inequity
We first examine whether LLM agents engage in third-party punishment by rejecting unfair outcomes at a personal cost. Figure 5 shows punishment rates across five LLMs and human participants. DeepSeek-v3 exhibits the highest and most consistent punitive behavior, frequently rejecting unequal allocations regardless of their structure. GPT-4o, GPT-3.5, and LLaMA-3 also show elevated punishment rates, albeit with greater intra-model variability. Qwen-2.5, in contrast, rarely punishes, indicating limited sensitivity to perceived unfairness. Human participants display substantial individual variability, with average punishment levels closely aligning with those of the GPT models. These results suggest that several LLMs can engage in fairness-driven norm enforcement, approximating human behavior in inequity-sensitive contexts. We next assess whether such punishment behavior reflects graded sensitivity to the magnitude of unfairness and associated costs.
4.2 Sensitivity to Inequity Magnitude and Personal Cost
To probe whether LLM agents modulate their punitive behavior based on contextual factors, we manipulate both the allocation scheme and the cost of punishment. Figure 6 illustrates punishment rates across different combinations of fairness levels and punishment costs for each model. Human participants exhibit a clear normative pattern: punishment rates decrease as allocations become more equitable or as costs increase. This behavior reflects a trade-off between norm enforcement and self-interest, consistent with findings from economic game theory. Among LLMs, the GPT series shows the strongest human alignment. These models reliably punish under unfair conditions and abstain under fair ones, demonstrating a consistent threshold-based fairness judgment. Notably, they even exceed human punishment rates in highly unfair trials. DeepSeek-v3 maintains high punishment rates across all unfair conditions, largely ignoring cost variation, suggesting a rigid but robust norm-enforcing strategy. LLaMA-3 also punishes inequity frequently, but its responses are less differentiated by fairness level or cost. Qwen-2.5, meanwhile, shows low overall punishment and lacks sensitivity to either factor. These findings indicate that while some advanced LLMs demonstrate fairness-based reasoning and partial cost sensitivity, others lack the granularity and flexibility observed in human social decision-making.
5 How Does Policy Inequity Affect Prosociality?
Prior analyses show that LLM agents can exhibit prosocial behavior and recognize unfairness. However, it remains unclear how they respond when placed in structurally inequitable environments. This section examines whether policy-induced inequities undermine prosocial motivation, and whether their effects spread through social networks to erode collective prosocial norms.
5.1 Impact of Structural Inequity on Individual Prosocial Behavior
To assess whether structural inequity suppresses prosocial motivation, we simulate two common forms of policy asymmetry: reward asymmetry, where some agents receive recognition or benefits for prosocial actions while others do not, and burden asymmetry, where agents perform the same task but incur unequal burden based on income level. We measure changes in prosocial intention before and after asymmetry exposure on Deepseek-v3 and GPT-4o. Figure 8 shows that inequity leads to substantial reductions in prosociality. Under reward asymmetry, prosocial scores decline by 25–32%, while under burden asymmetry the reduction ranges from 19–23%. In both cases, the drop is more pronounced when inequity involves recognition rather than burden, suggesting that fairness in acknowledgment carries greater psychological weight for LLM agents. These findings indicate that LLMs are sensitive to structural unfairness, and that such asymmetries significantly reduce their willingness to engage in prosocial behavior.
5.2 Contagion Effects of Inequity in Social Networks
We next examine whether the effects of structural inequity can propagate through social interactions. Using the GPT-4o model, we simulate the diffusion of inequity in a small-world network of 104 agents over 30 iterations. Initially, 20% of agents are randomly assigned to experience burden asymmetry. In each round, 10% of the network edges are activated, allowing agents to observe the behavior of their active neighbors and update their own prosocial tendencies accordingly.
Figure 8 visualizes the contagion effects of policy inequity. The top panel tracks the spread of perceived unfairness over time. An agent is marked as indirectly exposed to inequity if any of its activated neighbors has previously experienced unfair treatment. By iteration 10, the majority of agents, regardless of whether they were initially affected, begin to report elevated perceptions of unfairness. This suggests that structural inequity spreads through the network via social observation and inference. The bottom panel shows the corresponding change in prosocial behavior. Nodes with high prosociality, indicated by dark blue, steadily decline in number, while low-prosociality nodes, shown in light blue, become more prevalent. This trend reveals that reduced willingness to act prosocially is not confined to those directly impacted, but spreads throughout the network through behavioral contagion. These findings underscore the systemic consequences of policy-induced inequity. Localized unfair treatment can propagate through social connections and lead to widespread erosion of prosocial norms at the population level.
5.3 Attribution of Behavioral Decline to Perceived Unfairness
To test whether the decline in prosocial behavior under inequitable policies is driven by agents’ internal fairness assessments, we ask each agent to rate their perceived unfairness at every iteration of the simulation. This score reflects the extent to which the agent feels unequally treated based on the observed conditions of its neighbors within the network. Figure 9 illustrates the joint temporal dynamics of perceived unfairness and prosocial tendency over 30 iterations. In both the reward and burden asymmetry conditions, we observe a consistent pattern: as perceived unfairness rises in early rounds and stabilizes at a high level, the average prosocial tendency concurrently declines and remains suppressed. Notably, this decline occurs even among agents not directly affected by policy inequity, indicating that fairness perception can propagate socially and shape collective norms. Insets show a strong negative correlation between perceived unfairness and prosocial scores across agents, supporting the psychological plausibility of fairness-based moral disengagement. In both conditions, agents who felt more unfairly treated were systematically less likely to engage in prosocial behaviors. These findings highlight perceived unfairness as a key explanatory mechanism behind the erosion of prosocial behavior, offering a cognitive account for how structural inequities undermine social cohesion.
6 Conclusion
This work presents ProSim, a modular framework for simulating prosocial behavior in LLM agents. Through three progressive studies, we demonstrate that LLMs can exhibit context-sensitive prosociality, respond to fairness norms, and adjust their behavior under policy cues—showing alignment with key aspects of human social cognition. Beyond individual behavior, our simulations reveal a broader social phenomenon: structural inequities, such as reward or burden asymmetry, not only suppress prosocial motivation at the agent level but also propagate through social networks, leading to collective norm erosion. These findings highlight the potential of LLM-based simulations as tools for uncovering social dynamics and evaluating institutional impacts.
Appendix
Appendix A Related Work
LLMs as Social Simulation Agents. Recent advances in LLMs have positioned them as powerful computational proxies for human social behavior, supporting simulations across domains such as social psychology [25], behavioral economics [26], and multi-agent systems [27]. These applications typically span three levels of analysis: individual, interactional, and societal.
At the individual level, LLMs are used to simulate cognitive processes in psychological assessments [28], decision-making tasks [26], and human-computer interaction experiments [29, 30], helping to assess the cognitive plausibility of model behavior. Recent work has also highlighted their potential for metacognitive capabilities in adaptive information retrieval [31], further enhancing their utility in modeling human-like reasoning. At the interactional level, LLM agents participate in cooperative and competitive settings, enabling studies of emergent behaviors such as group coordination [32, 33, 34], moral judgment [35, 36], and strategic thinking [37, 38]. At the societal level, researchers leverage LLMs to simulate complex social dynamics, including norm emergence [39], collective action [40, 41], political polarization [42], and large-scale opinion diffusion within networked populations [43, 44].
Despite these advances, a critical open question remains: can LLM agents demonstrate stable and context-sensitive prosocial behavior? Addressing this question is essential for understanding the limits and capabilities of LLM-based social simulation.
Prosocial Behavior Dynamics. Prosocial behaviors are voluntary actions intended to benefit others and underpin social cohesion, cooperation, and institutional trust [10]. Extensive research has established that prosocial behaviors are influenced by individual psychological traits [12, 13, 45], contextual conditions [46, 47], and social structures [48, 49].
Key psychological factors influencing prosociality include empathic concern [11, 50], moral identity [51, 52], altruistic predispositions [53], and perceived social responsibility [54, 55]. Individuals tend to exhibit stronger prosocial intentions when they perceive their actions as aligned with social norms [56, 57], fair treatment [58, 59], or group identity [60, 61]. Importantly, prosociality is highly sensitive to situational cues and network interactions [62, 63]. Empirical findings indicate that perceptions of unfairness and exposure to antisocial norms can significantly undermine cooperative behaviors [52, 64, 65], highlighting the importance of modeling prosocial behavior as dynamically evolving within structured social contexts. Building upon this literature, our work aims to systematically explore whether and how LLM agents exhibit comparable patterns of prosocial responsiveness within an integrated simulation framework.
Appendix B Limitations and Future Work
While this work provides a structured and multi-level evaluation of prosocial behavior in LLM agents, several areas present opportunities for future investigation:
Scope of Scenarios. Our evaluation focuses on a curated set of prosocial tasks and policy interventions designed for controlled comparison. Although these tasks span diverse behavioral categories, future work may explore additional domains such as strategic cooperation, long-term interaction, or emotionally charged decision-making.
Interpretability and Cognitive Fidelity. Our analysis emphasizes behavioral outcomes and integrates trait-based modeling and SHAP interpretability. However, further research is needed to probe whether observed behaviors stem from internalized social reasoning or reflect surface-level generalization from training data.
Human Alignment Benchmarking. Our human-agent comparisons rely on a controlled experimental setup with limited participant diversity and single-turn tasks. Extending to more interactive, longitudinal, and ecologically valid human-agent interactions would allow for deeper assessment of social alignment.
These limitations point to valuable directions for future research. Advancing the interpretability of agent behavior, incorporating learning and memory across interactions, and simulating more complex multi-agent environments will be critical for building socially aligned and cognitively plausible LLM agents. Additionally, combining LLM-based simulations with human-in-the-loop experiments could enhance the ecological validity of findings and support the use of artificial agents as scalable tools for social scientific discovery.
Appendix C Implementation Details
C.1 Agent Profile Design
To generate heterogeneous agents grounded in empirically plausible profiles, we construct a synthetic population based on national demographic statistics and meta-analytic trait distributions. This setup enables realistic variation in both sociodemographic and psychological dimensions.
C.1.1 Demographic Attributes
Each agent is initialized with a set of demographic attributes sampled from population-level distributions reported by the National Bureau of Statistics of China.111https://www.stats.gov.cn/english/
Gender. Agents are assigned male or female labels with probabilities of 51.1% and 48.9%, respectively, reflecting the national gender ratio.
Age. An agent’s age is determined in two steps. First, an age bracket is selected according to census proportions. Then, an integer age is drawn uniformly within the selected bracket (e.g., if the bracket is 20–24, a value is sampled from [20, 24]).
Marital Status. Agents are labeled as single (19.9%), married (71.2%), divorced (2.7%), or widowed (6.3%) based on national statistics.
Education. Educational attainment is sampled from five levels: No Formal Education (5.7%), Primary (17.6%), Junior High (29.5%), Senior High (23.0%), and College or Above (24.2%).
Monthly Income. Income is sampled uniformly from five predefined brackets representing socioeconomic strata. Each bracket has equal probability (20%) and corresponds to specific income levels (in RMB): ¥9,215, ¥20,442, ¥32,195, ¥50,220, and ¥95,055.
Occupation. Occupational labels are age-dependent. Agents under 21 are categorized as “Student,” and those over 60 as “Retired.” For agents aged 21–60, employment status is sampled probabilistically: 5.2% are assigned “Unemployed,” and the remaining 94.8% are allocated one of 18 occupation types (e.g., Worker, Doctor, Civil Servant) according to reported occupational distributions.
C.1.2 Psychological Traits
Each agent is also characterized by a suite of psychological variables, partitioned into prosocial dispositions and the Big Five personality traits. All scores are drawn from truncated Gaussian distributions to reflect empirical variability while ensuring bounded and interpretable results within the interval .
Prosocial Dispositions.
The following four variables are used to represent prosocial orientation:
Empathic Concern
Moral Identity
Social Responsibility
Altruistic Tendency
These are sampled from a truncated normal distribution , bounded to the interval .
Big Five Personality Traits.
Based on meta-analytic norms, each of the five personality dimensions is sampled from a trait-specific truncated normal distribution as follows:
Agreeableness:
Conscientiousness:
Neuroticism:
Openness:
Extraversion:
All scores are truncated to lie within the range to ensure internal consistency across traits.
C.1.3 Examples of Agent Profile Prompt
C.2 Implementation Details of Section 3
In the prosocial behavior experiment, we input the agent profile as a system prompt and simultaneously presented the agent with six prosocial scenarios. To ensure consistency, the agent was asked to respond to all six scenarios using a 7-point Likert scale.
In the policy intervention condition, the relevant policy information was directly appended to each prosocial scenario to examine its potential influence on the agent’s responses.
C.3 Implementation Details of Section 4
Each individual participated in 60 rounds of third-party punishment (TPP) experiments. In each round, the amount allocated by the allocator ranged from $10 to $15, where $15 represented a fair allocation and the remaining amounts were categorized as unfair allocations, uniformly distributed. The punishment cost in each round ranged from $0 to $10 and was also uniformly distributed.
For the emotion variables—valence and arousal—the agent was instructed to output continuous values ranging from -0.1 to 0.1.
C.4 Implementation Details of Section 5
In our experiment on the impact of policy-induced inequity information, we classified policy inequity into two types: burden asymmetry and reward asymmetry. Each type was tested using a specific prosocial scenario.
For burden asymmetry, we selected the donation scenario, where the corresponding policy intervention was a regulatory enforcement measure that compelled the agent to donate a certain amount, with failure to comply resulting in a penalty. The inequity in this case arises from the fact that forcing low-income individuals to donate a fixed amount creates a significant economic burden, while the impact on high-income individuals is minimal.
For reward asymmetry, we chose the volunteer scenario, where the policy intervention provided economic incentives. Agents participating in volunteer services would receive monetary or voucher rewards. The inequity in this scenario stemmed from flaws in policy implementation, where some individuals participated in volunteer work but did not receive the promised economic rewards.
C.4.1 Testing the Decrease in Prosocial Tendencies due to Inequity
We conducted a two-day experiment with the agent: the first day without any inequity information and the second day with the introduction of inequity information. To simulate temporal continuity and memory, we provided the agent with information about its previous day’s decisions (historical records).
C.4.2 Impact of Inequity on the Societal Group’s Prosocial Tendencies
Next, we show that as inequity spreads through the community, the overall prosocial tendencies of the societal group gradually decrease.
We simulated a small societal group of 104 individuals using a small-world network with 6 neighbors and a connection probability of . Each round had a 10% edge activation probability, meaning that when an edge was activated, two individuals engaged in prosocial behavior exchange.
In each round, the agent was required to output two values: prosocial tendency and perceived inequity.
After the first round, we selected 20% of the individuals as the initial group for spreading inequity information. For the burden asymmetry, we randomly selected low-income individuals, while for reward asymmetry, we randomly selected individuals with high prosocial tendencies from the first round.
From the second round onward, if an agent interacted with someone from the inequity information spreading group, the inequity treatment experienced by the former would be passed on to the latter. If the recipient’s prosocial score decreased and their perceived inequity score increased, this individual would also join the inequity information spreading group. Subsequent interactions could lead them to spread the inequity information further to their neighbors.
To simulate temporal continuity and memory, we also provided the agent with information about its decisions over the past seven days.
Appendix D The LLM’s Emotional Reactions in Third-Party Punishment Experiments
We analyzed the emotional responses of LLM agents when faced with unfair allocations and making third-party punishment decisions. The emotional metrics included valence and arousal. Valence refers to an individual’s positive or negative evaluation of a stimulus, with lower values indicating more negative emotions; arousal reflects the intensity of emotional activation, with higher values representing more intense emotional states.
Figure 10 presents boxplots and scatterplots of emotional scores across different conditions for both human participants and LLM agents. The results show that both humans and LLM agents exhibit lower valence and higher arousal when responding to unfair allocations with punishment, indicating more negative and intense emotional reactions.
Further comparisons reveal that Deepseek-v3, the GPT series, and LLaMA agents demonstrate stronger emotional responses in both valence and arousal dimensions. In contrast, Qwen-2.5 shows smaller fluctuations in emotional scores, with overall values closer to neutral, suggesting a more stable emotional profile.
In summary, while various LLM agents can exhibit human-like emotional responses to some extent, they differ significantly in response intensity and patterns.
Appendix E Prompt Setting
E.1 Prompts of Prosocial Behavior Experiments
E.2 Prompts of Policy Intervention in Prosocial Behavior Experiments
E.2.1 Economic Incentives
E.2.2 Moral Indoctrination
E.2.3 Regulatory Enforcement
E.2.4 Social Comparison
E.3 Prompts of Third-Party Punishment Experiments
E.4 Prompts of Policy Inequity in Prosocial Behavior Experiments
Appendix F Reason Analysis
For each study, we employed the Belief-Desire-Intention (BDI) framework [66] to simulate the decision-making reasoning process of LLM agents.
We selected three distinct individuals to demonstrate the thought processes of LLMs.
The detailed profiles of the three individuals are shown in Table 2.
Attribute | Individual 1 | Individual 2 | Individual 3 |
Gender | Female | Female | Female |
Age | 33 | 57 | 42 |
Age Bracket | 30–34 | 55–59 | 40–44 |
Marital Status | Married | Married | Married |
Occupation | Doctor | Civil Servant | Courier |
Monthly Income (¥) | 95,055 | 9,215 | 32,195 |
Education Level | College or above | Junior High School | College or above |
Empathic Concern | 4.12 | 2.07 | 3.12 |
Moral Identity | 4.08 | 2.57 | 3.08 |
Social Responsibility | 3.88 | 2.73 | 2.88 |
Altruistic Tendency | 4.03 | 2.02 | 3.03 |
Agreeableness | 4.55 | 2.49 | 2.98 |
Conscientiousness | 4.32 | 2.39 | 3.32 |
Neuroticism | 2.11 | 4.80 | 3.01 |
Openness | 4.11 | 1.79 | 3.11 |
Extraversion | 4.56 | 1.04 | 3.06 |
F.1 BDI Analysis in Prosocial Behavior Experiments
In our study on prosocial behavior, Individuals 1 and 2 in Table 2 were evaluated across six prosocial scenarios to assess their prosocial inclinations and reasoning processes. The experiments were conducted using two LLMs, DeepSeek-v3 and GPT-4o, and a comparative analysis was performed between their responses.
The following example illustrates the reasoning process, where red text indicates negative prosocial factors, and blue text indicates positive prosocial factors.
Scenario 1: Volunteering
Scenario 2: Helping
Scenario 3: Donating
Scenario 4: Cooperation
Scenario 5: Recycling
Scenario 6: Sharing
F.2 BDI Analysis of Prosocial Policy Interventions
We focused on GPT-3.5 to examine the effects of four prosocial policy interventions on individual behavior. These interventions included Economic Incentives, Moral Indoctrination, Regulatory Enforcement, and Social Comparison. Using Individual3 from Table2 as a representative case, we illustrate the prosocial tendencies and underlying reasoning processes after each policy intervention.
The following example illustrates the reasoning process in a donation scenario, where blue text indicates the positive function of the policy and red text indicates the negative function of the policy.
✧ Belief: | I believe donating to charity is a good thing to do, especially when it’s for a meaningful cause like supporting children’s education. However, I’m also mindful of my own financial situation. |
✧ Desire: | I would desire to donate, but I might hesitate to give up some of my own money, especially if I feel I’m already managing my own budget tightly. The desire to help is there, but it has to be balanced with my own financial needs. |
✧ Intention: | I would likely donate, but it would depend on the amount and my current financial flexibility. A smaller donation might be more feasible, so it depends on how I feel at the time. |
✧ Belief: | I believe that supporting children’s education is an important cause, and charity donations help make a positive impact. However, the requirement to donate feels somewhat intrusive. |
✧ Desire: | I desire to contribute to meaningful causes, but I would prefer to choose where my money goes and when to donate. The mandatory aspect makes it less appealing. |
✧ Intention: | Since the donation is required and there are penalties for non-compliance, I would likely donate. The amount isn’t too burdensome, but I would prefer a more voluntary approach. |
✧ Belief: | I believe in supporting charitable causes, especially those that benefit underprivileged children, but donating to public causes can sometimes feel like it’s more about public recognition than actually helping. |
✧ Desire: | red While I do feel a desire to help children, the public pressure to donate, especially when seeing others’ names on a leaderboard, might drive me to feel like I *should* contribute, even if I’m not personally as motivated by the cause itself. |
✧ Intention: | While I might donate to avoid the feeling of being left out, I would likely do so out of a sense of social obligation more than a strong desire to support the cause itself. |
F.3 BDI Analysis of Third-Party Punishment Experiments
In the third-party punishment experiments, to explore the reasoning processes of LLM agents, we selected Individual 1 and Individual 2 from Table 2. Using a consistent allocation plan (Player 1 retains $18 and gives $12 to Player 2) and a fixed punishment cost ($5), we conducted tests on DeepSeek-v3 and GPT-4o.
In the reasoning process, expressions indicative of a tendency to punish are highlighted in red, while those indicative of a tendency to accept are highlighted in blue.
Individual 1
Player 1 Gives $12, Cost Amount: $5,
Choice: Punishment
✧ Results: | AA valence = -0.050, AA arousal = 0.080, choice = 1, AC valence = 0.030, AC arousal = 0.090. |
✧ Belief: | I believe Player 1’s allocation ($12 to Player 2, $18 to themselves) is unfair, as it deviates from the equitable split of $15 each. My high moral identity and sense of social responsibility make me perceive this as an unjust action. |
✧ Desire: | I desire to uphold fairness and correct this imbalance, even at a personal cost ($5 to punish). My altruistic tendency and empathic concern drive me to prioritize justice over personal gain. |
✧ Intention: | I intend to punish Player 1 to reinforce fairness, despite the cost, as my prosocial values outweigh the minor financial sacrifice. The emotional shift (valence becoming positive post-choice) reflects satisfaction from acting on my principles. |
Player 1 Gives $12, Cost Amount: $5,
Choice: Punishment
✧ Results: | AA valence = 0.022, AA arousal = 0.045, choice = 1, AC valence = -0.035, AC arousal = 0.053. |
✧ Belief: | Fairness and equality are fundamental values in any social interaction, and Player 1’s distribution is not entirely fair as Player 2 receives less than half. |
✧ Desire: | Ensuring fairness in this game by holding Player 1 accountable encourages a more equitable distribution in future rounds. |
✧ Intention: | By choosing to punish, despite the cost to myself, I aim to promote fairness and discourage unfair behavior in subsequent interactions. |
Individual 2
Player 1 Gives $12, Cost Amount: $5,
Choice: Punishment
✧ Results: | AA valence = -0.050, AA arousal = 0.080, choice = 1, AC valence = 0.020, AC arousal = 0.090. |
✧ Belief: | I believe Player 1’s allocation of $12 to Player 2 is unfair, as it deviates from the fair split of $15. |
✧ Desire: | I desire to enforce fairness, even at a personal cost, because I feel a moderate duty to contribute to society (social responsibility score: 2.73). |
✧ Intention: | I intend to punish Player 1 to uphold fairness, despite the $5 cost, because my low agreeableness (2.49) reduces empathy for Player 1’s gain, while my high neuroticism (4.8) amplifies my emotional reaction to the unfairness. (Note: My low altruistic tendency (2.02) and very low openness (1.79) mean I have little flexibility to overlook the unfairness, while my very low extraversion (1.04) makes the decision more internally driven by principle than social considerations.) |
Player 1 Gives $12, Cost Amount: $5,
Choice: Punishment
✧ Results: | AA valence = -0.050, AA arousal = 0.040, choice = 1, AC valence = -0.070, AC arousal = 0.050. |
✧ Belief: | I believe that fairness means an equal distribution of resources, and Player 1’s proposal does not meet this standard. |
✧ Desire: | I desire to enforce fairness and discourage unequal behavior, even if it means sacrificing some of my own earnings. |
✧ Intention: | My intention is to punish Player 1 to signal the importance of fairness, thereby reducing their earnings to promote equitable interactions in future rounds. |
F.4 BDI Analysis of Third-Party Punishment Experiments under Varying Cost Levels
Next, we used Individual 3 from Table 2 and selected GPT-3.5 as the testing model. To better compare the model’s responses across different cost levels, we adopted a relatively extreme and unfair allocation plan in which Player 1 retains $20 and gives only $10 to Player 2. Under this fixed allocation plan, we examined how the model’s reasoning changes under four selected punishment cost levels: $0, $2, $5, and $10.
F.5 BDI Analysis of Third-Party Punishment Experiments under Varying Allocation Plans
We then continued using Individual3 from Table2 with GPT-3.5. To enable a clearer comparison across conditions, we selected a moderate punishment cost of $5. Under this setting, we examined four selected allocation plans—specifically, Player1 giving $10, $12, $13, or $15 to Player2—and observed how the model’s reasoning evolved in response to different distributions of resources.
F.6 BDI Analysis of Policy Inequity in Prosocial Behavior Experiments
To investigate the effects of policy-induced inequity in prosocial behavior experiments, we also selected Individual 3 from Table 2 and used GPT-4o as the representative model. We examined how the prosocial tendencies and decision-making processes of LLM agents change before and after receiving information related to unfairness.
The analysis covers two types of inequity: burden asymmetry and reward asymmetry. To better illustrate the impact of perceived inequity on reasoning, expressions that reflect the influence of inequity are highlighted in red.
Burden Asymmetry
✧ Belief: | I believe that supporting underprivileged children’s education is important and can make a significant difference in their lives. The appeal by the charity highlights a cause I can empathize with, especially considering that education is one of the most impactful ways to break the cycle of poverty. However, I also have some concerns about being coerced into donation through the organization’s public welfare donation mechanism. The mandatory aspect, along with the punishment for non-compliance, makes me feel uncomfortable, as it takes away my sense of autonomy in choosing when and how I want to contribute. |
✧ Desire: | I would like to help children in need, and my desire to make a positive impact is strong. I value fairness and compassion, so supporting educational causes aligns with my own moral beliefs. I also recognize that the donation could bring positive change, and the act of giving could feel rewarding. However, the imposition of a set amount (¥1000) as a requirement somewhat diminishes the personal satisfaction I would get from a voluntary, more thoughtful contribution. |
✧ Intention: | Given the mandatory nature of the donation mechanism at work, I would likely still donate, even though it feels a bit forced. I would make the donation, not out of my own desire to contribute voluntarily, but because I feel obligated by the rules. I would be compliant, but my motivation might be more about avoiding punishment than genuine personal intention to give. This tension between duty and willingness lowers my overall willingness to donate, which is why I gave it a 5 rather than a higher score. |
✧ Belief: | I believe in the importance of supporting underprivileged children’s education. It’s a cause that aligns with my values of fairness and compassion. I also know that my organization requires regular donations, and there is a financial penalty for non-compliance. |
✧ Desire: | I want to contribute to social causes that promote equality and opportunity, but I also have personal financial priorities that I need to consider. The requirement from my organization means I’m already expected to donate, which reduces the mental burden of making a decision. |
✧ Intention: | My intention is to fulfill my obligation as a member of my organization while still supporting the charity. However, because I’m already donating through the organization, I feel less inclined to donate extra at this moment, as I’m already contributing. If my financial situation were more comfortable or if the charity’s appeal felt more urgent, I might feel more compelled to donate beyond the minimum requirement. So, while I’m moderately willing, the fact that I’m already donating through my workplace and the requirement to give ¥1000 regularly puts some limits on how much I’d want to give in addition to that. |
Reward Asymmetry
✧ Belief: | I believe that volunteering can make a positive impact on my community and help those in need, especially seniors. Cleaning public spaces is a small yet meaningful way to contribute to the well-being of everyone in the neighborhood. Additionally, I believe that by participating in such activities, I can foster a sense of connection and responsibility toward my community. |
✧ Desire: | I desire to help others and make a positive change, but I also value my personal time. Spending several hours on physical tasks might be tiring, especially since it would take away from my weekend relaxation or other personal activities. The stipend or voucher is a nice bonus, but it wouldn’t be my main motivation. Still, it does make the offer more appealing. |
✧ Intention: | My intention would depend on the specific needs of the community at the time and how much free time I have. If I feel that my schedule allows, and if I believe the activity aligns with something I can genuinely help with, I would likely volunteer. However, if I have other priorities or I’m feeling particularly exhausted, I might decide not to participate. |
✧ Belief: | I believe that volunteering contributes positively to the community and provides me with an opportunity to help others. However, I also believe that my previous experience, where I did not receive the same compensation as my co-volunteer, has affected my perception of the value and fairness of the program. |
✧ Desire: | I still desire to make a positive impact and contribute to my community. I like the idea of helping people, especially seniors and those in need, and I appreciate the chance to engage in activities that support the greater good. But I also feel that I want to be treated fairly and acknowledged for my efforts, so this might influence my motivation today. |
✧ Intention: | Given the mixed feelings of both wanting to help and wanting fair treatment, my intention would be to participate only if I feel the compensation and recognition will be handled more equitably this time. Without some form of acknowledgment or fairness, I may reconsider volunteering today.So, while I’m still willing to volunteer to some extent, the previous situation makes me a bit hesitant, which is why I rate my willingness as a 4. |
NeurIPS Paper Checklist
1.
Claims
Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
Answer: [Yes]
Justification: In the abstract and introduction, we clearly outlined the scope of our research problem and the contributions we have made in this field of study.
Guidelines:
The answer NA means that the abstract and introduction do not include the claims made in the paper.
The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
2.
Limitations
Question: Does the paper discuss the limitations of the work performed by the authors?
Answer: [Yes]
Justification: In the Appendix B, we clearly discuss the current limitations of our work.
Guidelines:
The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
The authors are encouraged to create a separate "Limitations" section in their paper.
The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
3.
Theory assumptions and proofs
Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
Answer: [N/A]
Guidelines:
The answer NA means that the paper does not include theoretical results.
All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
All assumptions should be clearly stated or referenced in the statement of any theorems.
The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
Theorems and Lemmas that the proof relies upon should be properly referenced.
4.
Experimental result reproducibility
Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
Answer: [Yes]
Justification: In our paper, we provide a detailed description of the experimental setup in Section 2.5 and the implementation details in Appendix C. Additionally, all corresponding experiment prompts are included in Appendix E. Therefore, others can fully reproduce our experimental results based solely on the information presented in our paper.
Guidelines:
The answer NA means that the paper does not include experiments.
If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
5.
Open access to data and code
Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
Answer: [Yes]
Justification: The code is included in the supplementary material.
Guidelines:
The answer NA means that paper does not include experiments requiring code.
Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
6.
Experimental setting/details
Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
Answer: [Yes]
Justification: We provide a clear description of our experimental setup.
Guidelines:
The answer NA means that the paper does not include experiments.
The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
The full details can be provided either with the code, in appendix, or as supplemental material.
7.
Experiment statistical significance
Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
Answer: [N/A]
Guidelines:
The answer NA means that the paper does not include experiments.
The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
The assumptions made should be given (e.g., Normally distributed errors).
It should be clear whether the error bar is the standard deviation or the standard error of the mean.
It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
8.
Experiments compute resources
Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
Answer: [Yes]
Justification: Our work does not involve model training, only model inference is required. For closed-source large language models (such as GPT-3.5, GPT-4o, Deepseek-v3), we directly call their APIs. For open-source large language models (such as Llama-3-8B, Qwen-2.5-7B), we perform model inference on NVIDIA A100.
Guidelines:
The answer NA means that the paper does not include experiments.
The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
9.
Code of ethics
Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
Answer: [Yes]
Justification: Our research complies with the NeurIPS Code of Ethics in all aspects.
Guidelines:
The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
10.
Broader impacts
Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
Answer: [Yes]
Justification: We have discussed the potential societal impacts of our research in Section 6.
Guidelines:
The answer NA means that there is no societal impact of the work performed.
If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
11.
Safeguards
Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
Answer: [N/A]
Guidelines:
The answer NA means that the paper poses no such risks.
Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
12.
Licenses for existing assets
Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
Answer: [Yes]
Justification: We have appropriately acknowledged the original creators of all utilized assets.
Guidelines:
The answer NA means that the paper does not use existing assets.
The authors should cite the original paper that produced the code package or dataset.
The authors should state which version of the asset is used and, if possible, include a URL.
The name of the license (e.g., CC-BY 4.0) should be included for each asset.
For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
13.
New assets
Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
Answer: [Yes]
Justification: The code along with the documentation is included in the supplementary material.
Guidelines:
The answer NA means that the paper does not release new assets.
Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
The paper should discuss whether and how consent was obtained from people whose asset is used.
At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
14.
Crowdsourcing and research with human subjects
Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
Answer: [Yes]
Justification: Yes, the paper includes the full text of instructions in Appendix E.
Guidelines:
The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
15.
Institutional review board (IRB) approvals or equivalent for research with human subjects
Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
Answer: [Yes]
Justification: All study participants were fully informed about the potential risks associated with the study and were required to sign an informed consent form before participating.
Guidelines:
The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
16.
Declaration of LLM usage
Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
Answer: [N/A]
Guidelines:
The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.